### v1

# VTAO-BiManip: Masked Visual-Tactile-Action Pre-training with Object Understanding for Bimanual Dexterous Manipulation

Abstract
Bimanual dexterous manipulation remains significant challenges in robotics due to the high DoFs of each hand and their coordination. Existing single-hand manipulation techniques often leverage human demonstrations to guide RL methods but fail to generalize to complex bimanual tasks involving multiple sub-skills. In this paper, we introduce VTAO-BiManip, a novel framework that combines visual-tactile-action pretraining with object understanding to facilitate curriculum RL to enable human-like bimanual manipulation. We improve prior learning by incorporating hand motion data, providing more effective guidance for dual-hand coordination than binary tactile feedback. Our pretraining model predicts future actions as well as object pose and size using masked multimodal inputs, facilitating cross-modal regularization. To address the multi-skill learning challenge, we introduce a two-stage curriculum RL approach to stabilize training. We evaluate our method on a bottle-cap unscrewing task, demonstrating its effectiveness in both simulated and real-world environments. Our approach achieves a success rate that surpasses existing visual-tactile pretraining methods by over 20%.
Figure 1: Left: examples of our bottle-cap turning dataset. Middle: simulation results of the bottle-cap task with Shadow Hand [1]. Right: the real-world deployment of the bottle-cap task with Leap Hand [2].
1College of Control Science and Engineering, Zhejiang University, Hangzhou, 310027, ChinaQi Ye (Corresponding author, qi.ye@zju.edu.cn) is with the College of Control Science and Engineering and the State Key Laboratory of Industrial Control Technology, Zhejiang University, and also with the Key Key Lab of CS&AUS of Zhejiang Province.This work was supported in part by the National Natural Science Foundation of China (Grant Number: 62088101, 62103372, 62233013).
I Introduction
Humans can perform various complex bimanual tasks in daily life while enabling robots to learn human-like bimanual manipulation skills is challenging due to the high degrees of freedom (DoFs) of each hand, and the coordination between hands. An algorithm needs to search in the high dimensional space for a two-hand configuration that conforms to physical laws and also meets the task requirements.
Solving a similar problem, existing work on single high DoFs robotic hand manipulation [3, 4, 5, 6] leverage priors from human demonstrations to tackle the challenge. In M2VTP [6], the prior knowledge of the human manipulation skill of bottle-cap turning is learned by a visual-tactile pre-training network and incorporated into a reinforcement learning (RL) method. However, the task in [6] involves turning a bottle cap with the bottle fixed on a table. In our task, the bottle is required to be grasped by the left hand and held in the air while the right hand performs bottle-cap turning as shown in Fig. 1 (middle). Considering the task similarity, the visual-tactile pre-training and RL method [6] is then applied to our task by collecting the human bimanual manipulation data and changing the state and action according to two hands.
However, the method failed in the RL training stage. The curves for the rewards and success rate stay near zero after thousands of iterations. The reasons can be attributed to two factors. Our task consists of three sub-tasks (grasp, hold, and bottle-cap turning). Learning the three skills and the coordination between these skills and two hands is required. 1) In M2VTP [6], the extra binary tactile modality helps the coordination between the fingers of a single hand in the bottle-cap turning by pretraining, but it cannot help the coordination between these skills and two hands as that much since touch status does not change in many segments. 2) When applying the RL method in [6] to our task, learning all these skills and coordination is treated as a black box and it is hard to learn all of them together.
To address the issues, in this paper, we propose VTAO-BiManip, a novel pretraining and RL learning framework for bimanual manipulation. It learns priors from more modalities during human manipulation and adopts a curriculum reinforcement learning strategy. Specifically, for the first issue, we observe that 1) Unlike binary tactile information, which only indicates physical contact, human hand actions offer richer and more intuitive priors on movement and coordination, effectively guiding downstream bimanual tasks; 2) the size and pose of objects determine how the hands must adjust and coordinate during manipulation, incorporating their perception is essential for achieving accurate and adaptable bimanual tasks. To learn these two priors, we add the action and object pose ’modality’ upon the visual-tactile pretraining network as shown in  Fig. 3. Notice that different from imitation learning which predicts the next action based on the input state, our action pretraining predicts successive future actions with masked input hand state. The masked and future action prediction strategy allows the pretraining network to benefit from a form of ’cross-modal regularization’ in learning the action prior due to the presence of other modality data. To pretrain our approach VTAO-BiManip, we use the same collection system of [6] and add a motion capture system to collect human visual-tactile manipulation data, hand movements, and object pose information.
For the second issue, a two-stage curriculum reinforcement learning framework is proposed: in the first stage, the bottles are fixed on the table, encouraging the left hand to grab the bottle and the right hand to unscrew the cap; in the second stage, the bottle is released, forcing the two-hand cooperative manipulation skill learning.
In summary, our contributions are as follows:
  • We propose VTAO-BiManip, a novel visual-tactile-action pretraining method with object understanding for bimanual dexterous manipulation. This approach leverages human data to pretrain models for action prediction and object state estimation.
  • We propose a curriculum reinforcement learning framework to address the challenge of multi-skill learning within a single task.
  • We pretrain VTAO-BiManip with collected human demonstrations and deploy to the downstream bimanual cap unscrewing task both in simulation and the real world.
II Related Works
II-A Multimodal Pretraining for Robotics
Pre-training methods utilizing unsupervised learning [7, 8, 9] have emerged as particularly effective in multimodal representation generalization. Due to the availability of extensive manipulation datasets that include visual and language modalities [10, 11, 12], some researches [13, 14, 15, 16] explores the representation pretraining of these modalities in robotic tasks, demonstrating their effectiveness and generalization capabilities.  [17] extends this by integrating action modalities with vision and language, enabling robots to predict and execute tasks. [18] improves downstream performance by predicting future motions of any point via visual pertaining. However, pretraining on action modalities has primarily focused on simple end-effectors, with limited exploration in dexterous manipulation. For tactile, studies such as [19, 20, 21, 6] highlight the effectiveness of visuo-tactile fusion in robotic manipulation. Yet, its application in bimanual tasks remains underexplored. To address this gap, we propose a novel approach, incorporating multimodal pretraining, including action prediction, into bimanual manipulation.
II-B Robotic Bimanual Manipulation
Learning-based bimanual manipulation has gained significant attention in robotics. While some studies have successfully utilized demonstration data [22, 23, 24], the lack of high-quality multi-fingered hand demonstrations often restricts their applicability to simple end-effectors. Efforts to collect such data for multi-fingered hands have been made [25, 26, 27], but challenges such as retargeting errors and latency limit their practical use.
On the other hand, sim-to-real approaches have also attracted growing interest in the domain of robotic manipulation [28, 29, 30, 31], demonstrating numerous successes across various tasks such as deformable object manipulation [32], open doors [33], and in-hand rotations [34]. However, most of these methods focus on either single dexterous hands [35, 36, 37, 38] or dual-gripper systems [39, 40]. Some progress has been made in dual-dexterous-hand manipulation [41, 42, 43].  [41] construct a dexterous bimanual manipulation platform, offering 20 manipulation tasks.  [42] employ multi-agent RL to train a bimanual handover task, incorporating a trajectory prediction model to bridge the sim-to-real gap.  [43] presents a system where robot hands are trained to play the piano using DRL. However, these works primarily rely on visual and proprioceptive modalities, which somewhat limit the performance of manipulation tasks. Multimodal data has proven effective in dual-gripper [44, 20, 45] and single dexterous hand tasks [46, 47, 48], its application in bimanual dexterous manipulation remains largely unexplored. Our work represents a pioneering effort in applying multimodal sensing to dual-dexterous-hand manipulation.
III Method
In this section, we first introduce our data collection system for capturing ego-centric human manipulation data, including Vision, Tactile, Action, and Object information (VTAO), detailed in Section III-A. Section III-B presents VTAO-BiManip, a novel pretraining framework that utilizes human manipulation priors for multimodal feature extraction and fusion. Finally, Section III-C outlines the development of a bimanual dexterous bottle-cap unscrewing task in simulation and the subsequent policy training using RL.
III-A VTAO Capture System for Human Bi-Manipulation
We develop a novel VTAO data collection system for capturing multimodal data during human bimanual manipulation, built upon our previous visual-tactile acquisition system [6].
III-A1 System Overview
The VTAO capture system includes: 1) tactile and motion acquisition gloves for both hands, 2) Hololens2  [^1] https://www.microsoft.com/en-us/hololens/buy for visual data capture, 3) the ChingMu motion capture (mocap) system  [^2] https://www.chingmu.com/ for recording hand movements and 6-DoF object pose, and 4) a personal computer responsible for data acquisition and alignment, as shown in Fig. 2. Each acquisition glove consists of two layers: an inner layer with 20 tactile sensors, as [6], and an outer layer with markers for motion capture. The mocap system comprises 16 cameras to capture the 6-DoF pose information of the bottle and the precise movements of the hand joints.
Figure 2: Our VTAO capture system for dual-hand human manipulation.
III-A2 Data Alignment and Unified Representation
Multimodal data are captured independently, with visual data sampled at 30 Hz, tactile data at 200 Hz, and mocap data at 1 kHz. Data from other modalities are synchronized with the visual modality using local timestamps, which are recorded by the personal computer during the acquisition. To unify coordinate systems across modalities, we place markers on the Hololens2 and calibrate the transformation between the camera and the rigid body coordinates of the Hololens2. Utilizing this transformation allows us to map the 6-DoF pose of the bottle, as well as the hand motion data, to the RGB camera coordinate system of the visual modality, facilitating model comprehension.
III-A3 Retarget Human Manipulation Trajectories
To unify bimanual action representations across pretraining and downstream tasks, we retarget human bimanual actions from the dataset onto the dual dexterous hands used in downstream tasks. Following DexRepNet [29], we regard retargeting as a nonlinear optimization problem. Our objective is to ensure similarity in distances from fingertips and finger segments to the wrist, in both robotic and human hands at all stages, resulting in natural robotic hand postures. Our objective function aligns with that presented in [25], formulated as:
$q_{t}^{\mathbf{R}*}=\underset{q_{t}^{\mathbf{R}}}{\operatorname{argmin}}\sum_{i=0}^{N}\|\mathbf{v}^{\mathbf{R}}_{i}\left(M^{\mathbf{R}},q_{t}^{\mathbf{R}}\right)-\mathbf{v}_{i}^{\mathbf{H}}\left(M^{\mathbf{H}},q_{t}^{\mathbf{H}}\right)\|^{2},$ (1)
where $\mathbf{v}_{i}^{\mathbf{R}}$ and $\mathbf{v}_{i}^{\mathbf{H}}$ signify the target vectors of the robotic and human hands respectively (from wrist to fingertips and wrist to finger segments). These are determined using forward kinematics of the dual Shadow Hand $M^{\mathbf{R}}$ and human hands $M^{\mathbf{H}}$, coupled with joint angles $q^{\mathbf{R}}_{t}$ and $q^{\mathbf{H}}_{t}$ at step $t$. $q_{t}^{\mathbf{R}*}$ denotes the optimized joint angles of the robotic hands at step $t$. We refine each step of the human bimanual manipulation trajectories within the dataset to generate retargeted robotic hand trajectories, which are subsequently utilized for the pre-training of the robots.
Figure 3: Overview of our VTAO-BiManip pre-train framework and its downstream application.
III-B Masked VTAO Transformer for Pretraining
To explore the benefits of incorporating multimodal human manipulation data into dual-hand robotic skill learning, we propose a novel pre-training framework, VTAO-BiManip, based on MAE [8]. This framework uses our collected VTAO data for pre-training. It extracts and integrates latent environmental representations to support the learning of downstream robotic manipulation tasks. VTAO-BiManip includes a VTAO encoder and a VTAO decoder, as shown in Fig. 3 (left).
III-B1 VTAO Encoder $E_{\theta}(V,C,A)\rightarrow(h_{\text{cls}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
The VTA encoder takes in the visual, tactile, and action pair (V, C, A) as input. The RGB image $V\in\mathbb{R}^{224\times 224\times 3}$, the binarized tactile signal from both hands $C\in\mathbb{R}^{40}$ (thresholded by $\lambda$), and the joint angles $A\in\mathbb{R}^{48}$, retargeted to the dual Shadow Hand, are processed via a Conv2D and two MLPs to form a patch sequence. Additionally, empty patches are added to align feature dimensions. These patches are used by the decoder to reconstruct the bottle information $\widehat{O}$. After adding positional encoding, the patch sequence becomes a token sequence. Following MAE [8], we randomly mask $r_{v}\in[0,1)$ of the visual tokens, $r_{c}\in[0,1)$ of the tactile tokens, $r_{a}\in[0,1)$ of the action tokens, and $r_{o}\in[0,1)$ of the empty tokens to form the input token sequence $(t_{v},t_{c},t_{a},t_{\text{null}})$. A CLS token $t_{\text{cls}}$ is inserted into this sequence as an aggregated representation for downstream tasks. The final token sequence $(t_{\text{cls}},t_{v},t_{c},t_{a},t_{\text{null}})$ is processed by a Transformer [24] encoder, producing a latent representation sequence $(h_{\text{cls}},h_{v},h_{c},h_{a},h_{o})$.
III-B2 VTAO Decoder $D_{\theta}(h_{v},h_{c},h_{a},h_{o},m)\rightarrow(\widehat{V},\widehat{C},\allowbreak[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}],\widehat{O})$
To enhance the feature extraction ability of the VTAO encoder $E_{\theta}$ for downstream tasks, we define four reconstruction objectives for the decoder: a) reconstruct the image $\widehat{V}$; b) reconstruct tactile signals $\widehat{C}$; c) reconstruct current actions and predict future actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$; d) reconstruct the rotation and size information of the bottle $\widehat{O}$, as shown in Fig. 3 (left). Here, $m$ denotes the mask tokens used for feature alignment.
Reconstruct Image $\widehat{V}$ and Tactile Signals $\widehat{C}$
Building on our previous work [6], we employ a Transformer Decoder and MLP layers to reconstruct the image $\widehat{V}$ and tactile signals $\widehat{C}$ from the corresponding tokens $h_{v}$, $h_{c}$, respectively. Serving as an auxiliary task, this approach has been demonstrated to benefit downstream applications.
Reconstruct Current Actions and Predict Future Actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$
To enable the encoder $E_{\theta}$ to acquire relevant prior knowledge of human manipulation tasks from single-frame VTAO input, we use a Transformer decoder and an MLP layer to reconstruct the dual-hand action of current step $\widehat{A}$ and the action sequences for the next p steps $[\widehat{A_{+1}},...,\widehat{A_{+p}}]$ from the corresponding action tokens $h_{a}$.
Reconstruct Bottle Informations $\widehat{O}$
Knowing the pose and size of the object is beneficial for downstream tasks such as bottle cap twisting. We enable the network to reconstruct the 6-DoF pose of the bottle relative to the camera coordinate system, along with the dimensions of the bottle cap and body, from object tokens $h_{\text{o}}$. This enables the encoder $E_{\theta}$ to better focus on the relevant information of the manipulated object.
Loss Design
We formulate the loss function as:
$L(\theta)=$ $W_{\text{img}}\cdot L(V,\widehat{V})+W_{\text{tac}}\cdot L(C,\widehat{C})+W_{\text{bot}}\cdot L(O,\widehat{O})$
$W_{\text{act}}\cdot L([A,A_{+1},...,A_{+p}],[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]),$ (2)
where $W_{\text{img}},W_{\text{tac}},W_{\text{bot}}$, and $W_{\text{act}}$ are weights and set to 1, 2, 5, 2, respectively. $L(\cdot)$ denotes the Euclidean distance between two vectors.
III-C VTAO RL for bimanual manipulations
III-C1 Problem Formulation
We model the bimanual manipulation task as a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$. Here, $\mathcal{S}$ and $\mathcal{A}$ represent the state space and action space. The policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$ maps states to actions. Transition dynamics are given by $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, and the reward function by $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. The discount factor $\gamma\in\left(0,1\right]$. Our objective is to maximize the expected discounted reward $J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}\gamma^{t}r\left(s_{t},a_{t}\right)\right]$ to train a policy network $\pi_{\theta}$. We employ the PPO algorithm to facilitate the learning of manipulation skills. Our RL architecture is illustrated in Fig. 3 (right).
III-C2 Task
To assess the effectiveness of VTAO pre-training on downstream tasks, we implement a bimanual bottle-cap unscrewing task in Isaac gym [49]. The setup positions the bottle body at the center of the table, with the left hand to its left in a pre-grasp gesture and the right hand fixed above it, as shown in the submitted video. The bottle body and cap are linked by a joint that rotates around the z-axis.
III-C3 State Space $\mathcal{S}$
We define the state space as $\mathcal{S}=\{V_{\text{obs}},C_{\text{obs}},\mathcal{P}\}$, where $V_{\text{obs}}$ is the ego-centric RGB image, $C_{\text{obs}}$ is the binary tactile signal from both hands, thresholded by $\lambda$, and $\mathcal{P}$ is proprioception data of both hands, including joint angles and velocities. To maintain feature representation from pre-training and improve stability in RL with sparse rewards, we freeze the parameters of the pre-trained VTAO model during RL training. The final feature vector input to the policy network is $\{h_{\text{cls}},\phi(\mathcal{P})\}$, where $h_{\text{cls}}=\mathrm{E}_{\theta_{f}}(V_{\text{obs}},C_{\text{obs}},A_{\text{obs}})$, with $\mathrm{E}_{\theta_{f}}$ representing the pre-trained VTAO encoder with frozen parameters. $A_{\text{obs}}\in\mathcal{P}$ represents the joint angles of both hands. $h_{\text{cls}}$ is the CLS token output by $\mathrm{E}_{\theta_{f}}$. $\phi(\cdot)$ is a linear layer.
III-C4 Action Space $\mathcal{A}$
In the simulation, we use Shadow Hand [1] as the manipulator. This five-fingered robotic hand features 24 DoFs. Excluding the thumb, the distal joints of the four fingers are tendon-driven, resulting in a total of 20 DoFs. We implement dual-hand manipulation, where the base link of the right hand is fixed and the left hand can move freely with six DoFs. Consequently, the action $a=\pi_{\theta}(s)\in\mathbb{R}^{46}$.
III-C5 Learning Curriculum
It is challenging to train precise dual-hand manipulation with non-fixed objects from scratch. To address this, we use a two-stage learning curriculum to stabilize the learning process. In the first stage, we fix the bottle body on the table, enabling the left hand to learn to grasp the bottle and the right hand to learn to unscrew the cap. In the second stage, we release the bottle body, allowing both hands to collaborate in unscrewing the cap.
III-C6 Reward Function $\mathcal{R}$
To coordinate the dual dexterous hands in complex action spaces across different training stages, our reward function $r_{n}(n\in\{1,2\})$ is divided into two phases. In each phase, the total reward is the sum of the individual rewards for the left and right hands, $r_{n}=r_{\text{left}_{n}}+r_{\text{right}_{n}}$. The reward function of the first stage $r_{1}$ is:
{$r_{\text{left}_{1}}&=r_{\text{hdis}}+r_{\text{fcon}}\\ r_{\text{right}_{1}}&=r_{\text{cang}}+r_{\text{cvel}}+r_{\text{fdis}}\end{aligned}\right..$ (3)
Here, $r_{\text{hdis}}=\alpha_{1}\cdot d_{\text{h2b}}$ is a distance penalty, where $d_{\text{h2b}}$ is the distance between the left hand and the bottle. $r_{\text{fcon}}=\alpha_{2}\cdot N_{\text{con}}$ incentivizes multiple contacts between the left hand and the bottle, with $N_{\text{con}}$ representing the number of contacts. The rotation angle reward, $r_{\text{cang}}=\alpha_{3}\cdot\min(a_{\text{c}},7.0)$, depends on the rotation angle of the bottle cap $a_{\text{c}}$. $r_{\text{cvel}}=\alpha_{4}\cdot C_{\text{flag}}\cdot v_{\text{c}}$ rewards rotation speed, where $v_{\text{c}}$ is the rotation speed of the cap, and $C_{\text{flag}}$ equals 1 if the right hand is in contact with the cap, and 0 otherwise. The term $r_{\text{fdis}}=\alpha_{5}\cdot\exp(-10\cdot d_{\text{fz}})$ encourages the right-hand fingertips to approach the bottle cap, with $d_{\text{fz}}$ being the z-axis distance between the fingertips and the height of the cap. The weights $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{5}$ are set to $-5$, $0.05$, $0.5$, $1.1$, and $0.5$, respectively.
The rewards for the second phase, $r_{2}$ are as follows:
{$r_{\text{left}_{2}}&=r_{\text{htdis}}+r_{\text{bdis}}+r_{\text{brot}}\\ r_{\text{right}_{2}}&=r_{\text{right}_{1}}\end{aligned}\right.,$ (4)
where $r_{\text{htdis}}=\beta_{1}\cdot\exp(-5\cdot d_{\text{h2t}})$ is the approaching reward, with $d_{\text{h2t}}$ representing the distance between the left hand and the target position. $r_{\text{bdis}}=\beta_{2}\cdot\exp(-10\cdot d_{\text{o2i}})$ encourages the bottle to remain near its initial position, with $d_{\text{o2i}}$ denoting the distance between the bottle’s current and initial position. The reward $r_{\text{brot}}=\beta_{3}\left(\frac{1}{|d_{\text{qua}}|+1}\right)$ encourages the axis of the bottle remains perpendicular to the table, where $d_{\text{qua}}=2\cdot\arcsin\left(\min\left(\lVert q_{\text{diff}}\rVert_{2},1.0\right)\right)$ , and $q_{\text{diff}}=q_{\text{bot}}\cdot q_{\text{ini}}$ represents the vector part of the quaternion difference between the bottle’s current quaternion $q_{\text{bot}}$ and its initial orientation $q_{\text{ini}}$. The weights $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are all set to 1.0.
IV Experiment Results
In this section, we design experiments to validate: 1) the effectiveness of multimodal fusion pre-training; 2) the efficacy of action prediction; and 3) the validity of object understanding.
IV-A Experiment Settings
IV-A1 Dataset
Using our VTAO acquisition system, we collect 216 human bimanual manipulation trajectories involving 26 different bottles. Each trajectory lasts 6 to 17 seconds and comprises a total of 61684 frames. Each frame contains aligned visual, tactile, and action data, along with bottle pose and size labels, as shown in the submitted video. Before training VTAO-BiManip, we preprocess the data by cropping images to $224\times 224$ pixels, binarizing tactile data with a $\lambda=0.4V$ threshold, aligning bottle pose with the camera coordinate system of the visual data, and retargeting human hand movements to the joint angles of the dual dexterous hands. The final data representations include images $V\in\mathbb{R}^{224\times 224\times 3}$, tactile data $C\in\mathbb{R}^{40}$, and action data $A\in\mathbb{R}^{48}$.
IV-A2 Metrics
We use success rate to evaluate the effectiveness of our method on downstream tasks. A task is considered successful if the dexterous hand rotates the bottle cap by more than half a turn. During training, we calculate the success rate as the average of the last ten manipulations in each parallel environment. For evaluation, we select the model from the 3500th iteration of the second training phase and measure the average success rate across 10 manipulations for each bottle in the evaluation set.
IV-A3 Implementation Details
We select 15 bottles from ShapeNet [50] with varying body and cap sizes, using 10 for training (Seen) and 5 for testing (Unseen). Our VTAO-BiManip runs on a system with an Intel Xeon Gold 6326 and an NVIDIA 3090. During pretraining, we use the AdamW optimizer with a learning rate of 2e-5. The mask ratios are set to $r_{v}=0.75$, $r_{c}=0.5$, and $r_{a}=0.5$. Training with a minibatch size of 8 takes about 2 hours for 300 epochs. For RL, we run 400 parallel environments and train using the PPO algorithm for roughly 4500 iterations, requiring about 62 hours to train the dual-hand manipulation strategy network.
IV-A4 Baselines
To validate the effectiveness of our method, we design several baselines with different pre-training structures, as shown in Table I. We designate our baselines in the form of VTAO. Here, V, T, and A correspond to the pretraining of visual, tactile, and action modalities, respectively, while O represents the reconstruction of bottle information. Both tactile and action modalities involve data from two hands. The action modality entails reconstructing the current frame and predicting the next five frames. The absence of a letter in the naming indicates that the corresponding modality is not used. For methods that do not involve pretraining, we denote them with the suffix ‘-Scr’.
TABLE I: Settings for baselines.
Methods Modality Pretrain JointPretrain ReconBottle
V v $\checkmark$ $\times$ $\times$
T t $\checkmark$ $\times$ $\times$
A a $\checkmark$ $\times$ $\times$
VT v+t $\checkmark$ $\checkmark$ $\times$
VTA v+t+a $\checkmark$ $\checkmark$ $\times$
VTO v+t $\checkmark$ $\checkmark$ $\checkmark$
VTA-Scr v+t+a $\times$ $\times$ $\times$
VTAO (Ours) v+t+a $\checkmark$ $\checkmark$ $\checkmark$
IV-B Effectiveness of pre-training with different modalities
* indicates that the method failed to train.
Figure 4: Qualitative results of different pre-train methods. The left is the training process; the right is the evaluation results
We design experiments to validate the effectiveness of our proposed multimodal joint pre-training. The experimental results are illustrated in the Fig. 4.
Compared to single-modality pretraining methods V, T, and A, as well as the multimodal method without pre-train VTA-Scr, the multimodal joint pretraining approach VTA demonstrates superior performance in downstream task manipulation success rates. This highlights the effectiveness of our multimodal joint pretraining strategy.
By comparing VT, VTA, VTO, and VTAO, we find that incorporating action modality prediction and object understanding into visuo-tactile fusion training enhances downstream task success rates. This validates the effectiveness and complementarity of the action prediction task and the object understanding task proposed in our approach.
Both Base and A, which only possess proprioception of the robotic hands without environmental perception, fail in training. In contrast, the pretraining method using only tactile perception, T, demonstrates a certain level of manipulation capability, effectively demonstrating the ability of tactile perception to gather environmental information and its effectiveness in enhancing operational performance.
IV-C Ablation Study
IV-C1 Effectiveness of Bimanual Pre-training
In our approach, we utilize tactile and action data from both hands for pre-training and downstream tasks. In this part, we validate the effectiveness of using both hands. We remove the left-hand action data in the VTA experiment (v1), resulting in version v2. Similarly, we remove the left-hand tactile data in the VT experiment (v3), producing version v4. As shown in Table II, using bimanual data consistently achieves higher success rates compared to using right-hand-only data, irrespective of the modality removed. This indicates that although the left hand mainly performs basic grasping and holding, its inclusion significantly enhances coordination and improves success rates.
TABLE II: Test success rates (%) on seen and unseen bottles across different experimental settings.
Methods Tac Act PredictAct ActToken Obj Seen Unseen
Ours dual dual 5steps $24\times 2$checkmark 63±4 51±7
v1 dual dual 5steps $24\times 2$$\times$ 54±23 39±15
v2 dual right 5steps 24 $\times$ 46±15 32±9
v3 dual $\times$ $\times$ $\times$ $\times$ 41±3 30±3
v4 right $\times$ $\times$ $\times$ $\times$ 37±17 20±7
v5 dual dual NoPredict $24\times 2$$\times$ 35±25 21±16
v6 dual dual 1steps $24\times 2$$\times$ 44±8 17± 5
v7 dual dual 5steps $1\times 2$$\times$ 31±4 18± 9
v8 dual dual 5steps $5\times 2$$\times$ 50±16 26± 8
Figure 5: Training process for the ablation study. Left: Bimanual manipulation validation; Middle: Action prediction horizon ablation; Right: Ablation of the number of action tokens.
IV-C2 Action Prediction Horizon
In our approach, we incorporate action prediction into the pre-trained model to forecast actions for the next 5 frames. In this part, we investigate the impact of the prediction horizon on downstream tasks. We evaluate the manipulation success rate of methods that either only reconstruct the action of the current frame (v5) or predict the action for the next frame (v6). The experimental results, as shown in Table II, indicate that compared to the baseline without action addition (v1), solely reconstructing the action of the current frame has a detrimental effect. However, when action prediction was included, the success rate improved. This highlights the importance of predicting future actions in pretraining for robotic policies.
IV-C3 Number of Action Tokens $\mathcal{N}$
Our method treats the action (angle values) of each joint in both hands as individual tokens, which are then fed into the Transformer. In this part, we explore different ways of segmenting the actions, considering hand actions based on each hand (v7) and each finger (v8). Experimental results, as shown in Table 2, demonstrate that setting the number of action tokens to 24×2 is more advantageous for VTA fusion.
IV-D Real World Experiment
Figure 6: Our manipulation platform.
We create a real-world experimental platform with dual Leap Hands [2] and a Unitree Z1 arm  [^3] https://www.unitree.com/z1, as shown in Fig. 6. The simulation task for bottle cap unscrewing is recreated in Isaac Gym [49] using dual Leap Hands. The final validation confirms that our method supports effective physical experimentation.
V Conclusion
This paper investigates the effectiveness of a pre-trained representation model integrating multi-modalities, action prediction, and object understanding for dexterous bimanual manipulation. We design a novel pre-training framework, VTAO-BiManip, which is pre-trained using human VTAO demonstrations. We integrate a learning curriculum into RL to address the challenge of multi-skill learning within a bimanual dexterous skill. The representation model is incorporated into the curriculum RL framework for bimanual manipulation. Experimental results show that our approach significantly outperforms baseline methods. This study represents a pioneering effort in multimodal bimanual manipulation. Future work may focus on: 1) developing large-scale multimodal bimanual datasets across diverse scenes and tasks; 2) exploring additional bimanual manipulation tasks; 3) leveraging unpaired existing 3D object labeling data.

### v2

@@
 Abstract
-Bimanual dexterous manipulation remains significant challenges in robotics due to the high DoFs of each hand and their coordination.
+Bimanual dexterous manipulation remains a significant challenge in robotics due to the high DoFs of each hand and their coordination.
 Existing single-hand manipulation techniques often leverage human demonstrations to guide RL methods but fail to generalize to complex bimanual tasks involving multiple sub-skills.
-In this paper, we introduce VTAO-BiManip, a novel framework that combines visual-tactile-action pretraining with object understanding to facilitate curriculum RL to enable human-like bimanual manipulation.
-We improve prior learning by incorporating hand motion data, providing more effective guidance for dual-hand coordination than binary tactile feedback.
+In this paper, we propose VTAO-BiManip, a novel framework that integrates visual-tactile-action pre-training with object understanding, aiming to enable human-like bimanual manipulation via curriculum reinforcement learning (RL).
+We improve prior learning by incorporating hand motion data, providing more effective guidance for dual-hand coordination.
 Our pretraining model predicts future actions as well as object pose and size using masked multimodal inputs, facilitating cross-modal regularization.
 To address the multi-skill learning challenge, we introduce a two-stage curriculum RL approach to stabilize training.
-We evaluate our method on a bottle-cap unscrewing task, demonstrating its effectiveness in both simulated and real-world environments.
+We evaluate our method on a bimanual bottle-cap twisting task, demonstrating its effectiveness in both simulated and real-world environments.
 Our approach achieves a success rate that surpasses existing visual-tactile pretraining methods by over 20%.
 Figure 1: Left: examples of our bottle-cap turning dataset.
-Middle: simulation results of the bottle-cap task with Shadow Hand [1].
-Right: the real-world deployment of the bottle-cap task with Leap Hand [2].
-1College of Control Science and Engineering, Zhejiang University, Hangzhou, 310027, ChinaQi Ye (Corresponding author, qi.ye@zju.edu.cn) is with the College of Control Science and Engineering and the State Key Laboratory of Industrial Control Technology, Zhejiang University, and also with the Key Key Lab of CS&AUS of Zhejiang Province.This work was supported in part by the National Natural Science Foundation of China (Grant Number: 62088101, 62103372, 62233013).
+Middle: simulation results of the bottle-cap task with Shadow Hand [shadowhand].
+Right: the real-world deployment of the bottle-cap task with Leap Hand [shaw2023leaphand].
+1College of Control Science and Engineering, Zhejiang University, Hangzhou, 310027, ChinaQi Ye (Corresponding author, qi.ye@zju.edu.cn) is with the College of Control Science and Engineering and the State Key Laboratory of Industrial Control Technology, Zhejiang University, and also with the Key Key Lab of CS&AUS of Zhejiang Province.This work was supported in part by NSFC under Grants (No.
+62088101, 62233013, 62293511) and Key Research and Development Program of Zhejiang Province (No.
+2025C01072).
 I Introduction
-Humans can perform various complex bimanual tasks in daily life while enabling robots to learn human-like bimanual manipulation skills is challenging due to the high degrees of freedom (DoFs) of each hand, and the coordination between hands.
-An algorithm needs to search in the high dimensional space for a two-hand configuration that conforms to physical laws and also meets the task requirements.
-Solving a similar problem, existing work on single high DoFs robotic hand manipulation [3, 4, 5, 6] leverage priors from human demonstrations to tackle the challenge.
-In M2VTP [6], the prior knowledge of the human manipulation skill of bottle-cap turning is learned by a visual-tactile pre-training network and incorporated into a reinforcement learning (RL) method.
-However, the task in [6] involves turning a bottle cap with the bottle fixed on a table.
-In our task, the bottle is required to be grasped by the left hand and held in the air while the right hand performs bottle-cap turning as shown in Fig.
-1 (middle).
-Considering the task similarity, the visual-tactile pre-training and RL method [6] is then applied to our task by collecting the human bimanual manipulation data and changing the state and action according to two hands.
-However, the method failed in the RL training stage.
-The curves for the rewards and success rate stay near zero after thousands of iterations.
-The reasons can be attributed to two factors.
-Our task consists of three sub-tasks (grasp, hold, and bottle-cap turning).
-Learning the three skills and the coordination between these skills and two hands is required.
-1) In M2VTP [6], the extra binary tactile modality helps the coordination between the fingers of a single hand in the bottle-cap turning by pretraining, but it cannot help the coordination between these skills and two hands as that much since touch status does not change in many segments.
-2) When applying the RL method in [6] to our task, learning all these skills and coordination is treated as a black box and it is hard to learn all of them together.
-To address the issues, in this paper, we propose VTAO-BiManip, a novel pretraining and RL learning framework for bimanual manipulation.
-It learns priors from more modalities during human manipulation and adopts a curriculum reinforcement learning strategy.
-Specifically, for the first issue, we observe that 1) Unlike binary tactile information, which only indicates physical contact, human hand actions offer richer and more intuitive priors on movement and coordination, effectively guiding downstream bimanual tasks; 2) the size and pose of objects determine how the hands must adjust and coordinate during manipulation, incorporating their perception is essential for achieving accurate and adaptable bimanual tasks.
-To learn these two priors, we add the action and object pose ’modality’ upon the visual-tactile pretraining network as shown in  Fig.
-3.
-Notice that different from imitation learning which predicts the next action based on the input state, our action pretraining predicts successive future actions with masked input hand state.
-The masked and future action prediction strategy allows the pretraining network to benefit from a form of ’cross-modal regularization’ in learning the action prior due to the presence of other modality data.
-To pretrain our approach VTAO-BiManip, we use the same collection system of [6] and add a motion capture system to collect human visual-tactile manipulation data, hand movements, and object pose information.
-For the second issue, a two-stage curriculum reinforcement learning framework is proposed: in the first stage, the bottles are fixed on the table, encouraging the left hand to grab the bottle and the right hand to unscrew the cap; in the second stage, the bottle is released, forcing the two-hand cooperative manipulation skill learning.
+Humans effortlessly perform diverse complex bimanual tasks like opening containers and assembling furniture in daily life via dynamic inter-hand coordination.
+Such capabilities rely on three fundamental components: 1) kinematic inter-hand coordination: dynamically adjusting joint-level movements between hands to maintain operational stability, 2) real-time object-state sensing: modifying manipulation strategies based on real-time state changes of manipulated objects, 3) temporally composition of complementary skill primitives: executing phase-dependent primitives (grasping, holding, and rotating) under physical constraints.
+While humans perform these tasks effortlessly, robotic dexterous bimanual manipulation remains an open challenge.
+The control complexity grows exponentially when coordinating two high-DoF hands (totaling 30+ joints) under physical constraints, particularly in dynamic scenarios with coupled dynamics between object states and manipulation requirements.
+Recent advances in dexterous manipulation [Rajeswaran2018DAPG, qin2022dexmv, wu2022ILAD, liu2024m2vtp] have demonstrated progress in single-hand scenarios by leveraging human demonstrations and multi-modal sensing.
+Particularly, Liu et al.
+[liu2024m2vtp] established a visual-tactile pre-training framework that enables transferable manipulation skills through masked autoencoding.
+However, these methods fundamentally rely on constrained task setups - for instance, fixing objects on tables to bypass bimanual coordination challenges.
+Directly extending such frameworks to dynamic bimanual tasks like bottle-cap twisting reveals three irreducible challenges: 1) kinematic coordination limitation: while vision and binary tactile sensing aids single-hand finger coordination [liu2024m2vtp], they fail to capture inter-hand kinematic synergy critical for dynamic bimanual manipulation.
+2) object-state decoupling: when the bottle tilts during cap rotation, the passive hand must respond based on real-time 3D pose changes.
+Current methods lack explicit perception of object pose and size, leading to incoherent hand movements under dynamic states.
+3) sub-skill composability: successful cap twisting requires temporally aligned sub-skills - the passive hand must establish stable grasping before the active hand initiates rotation.
+However, single-stage policies in prior work [liu2024m2vtp] lack mechanisms to coordinate these interdependent sub-processes.
+To address the issues, in this paper, we propose VTAO-BiManip, a novel framework extending [liu2024m2vtp] through two key innovations: visual-tactile-action pre-training method with object understanding for multi-modal perception and curriculum reinforcement learning for action control policy learning.
+To address kinematic coordination blindness and object-state decoupling, we introduce hand action and object-centric understanding as complementary modalities during pre-training: 1) while prior work relies on visual-tactile fusion, we introduce hand joint angles as explicit kinematic observables and predict future actions to model inter-hand synergy.
+This enables the policy to dynamically coordinate bimanual motions; for instance, progressively adjusting the finger joint angles of the passive hand to counteract active-hand torque during cap twisting.
+Notably, unlike imitation learning which predicts next-step actions from current states, our action pre-training predicts future action sequences using masked multi-modal inputs.
+2) we also augment the pre-training framework with object pose and size estimation, forcing the encoder to learn geometry-aware representations.
+By reconstructing the bottle’s 6D pose and 3D dimensions from visual-tactile-action inputs, the model gains explicit awareness of dynamic state changes.
+During deployment, this allows real-time adaptation.
+To address sub-skill composability, a two-stage curriculum reinforcement learning framework is proposed.
+In the first stage, the bottles are fixed on the table, encouraging the left hand to grab the bottle and the right hand to unscrew the cap.
+This phase reduces exploration complexity, building foundational sub-skills of both hands.
+In the second stage, the bottle is physically released, forcing the two-hand cooperative manipulation skill learning.
 In summary, our contributions are as follows:
-  • We propose VTAO-BiManip, a novel visual-tactile-action pretraining method with object understanding for bimanual dexterous manipulation.
-This approach leverages human data to pretrain models for action prediction and object state estimation.
+  • We propose VTAO-BiManip, a novel visual-tactile-action pre-training method with object understanding for bimanual dexterous manipulation.
+This approach leverages human data to pre-train models for action prediction and object state estimation.
   • We propose a curriculum reinforcement learning framework to address the challenge of multi-skill learning within a single task.
-  • We pretrain VTAO-BiManip with collected human demonstrations and deploy to the downstream bimanual cap unscrewing task both in simulation and the real world.
+  • We propose a bimanual bottle-cap twisting task where our VTAO-BiManip pre-training and curriculum RL framework succeeds both in simulation and the real world.
 II Related Works
 II-A Multimodal Pretraining for Robotics
-Pre-training methods utilizing unsupervised learning [7, 8, 9] have emerged as particularly effective in multimodal representation generalization.
-Due to the availability of extensive manipulation datasets that include visual and language modalities [10, 11, 12], some researches [13, 14, 15, 16] explores the representation pretraining of these modalities in robotic tasks, demonstrating their effectiveness and generalization capabilities.
-[17] extends this by integrating action modalities with vision and language, enabling robots to predict and execute tasks.
-[18] improves downstream performance by predicting future motions of any point via visual pertaining.
-However, pretraining on action modalities has primarily focused on simple end-effectors, with limited exploration in dexterous manipulation.
-For tactile, studies such as [19, 20, 21, 6] highlight the effectiveness of visuo-tactile fusion in robotic manipulation.
-Yet, its application in bimanual tasks remains underexplored.
-To address this gap, we propose a novel approach, incorporating multimodal pretraining, including action prediction, into bimanual manipulation.
+Pre-training methods that use unsupervised learning [radford2021clip, he2022mae, oord2018representation] have emerged as particularly effective in multimodal representation generalization.
+Due to the availability of extensive manipulation datasets that include visual and language modalities [goyal2017sth_sth_v2, grauman2022ego4d, Damen2022epic_kitchen], some researches [ma2023liv, nair2022r3m, karamcheti2023voltron, xiao2022mvp] explore the pretraining representation of these modalities in robotic tasks, demonstrating their effectiveness and generalization capabilities.
+[brohan2023rt2] extends this by integrating action modalities with vision and language, enabling robots to predict and execute tasks.
+[wen2023atm] improves downstream performance by predicting future motions of any point via visual pertaining.
+However, pre-training on action modalities has primarily focused on simple end-effectors, with limited exploration in dexterous manipulation.
+For tactile, studies such as [chen2023VTT, mejia2024hearing, gano2024vt, liu2024m2vtp] highlight the effectiveness of visuo-tactile fusion in robotic manipulation.
+However, its application in bimanual tasks remains underexplored.
+To address this gap, we propose a novel approach that incorporates multimodal pre-training, including action prediction, into bimanual manipulation.
 II-B Robotic Bimanual Manipulation
 Learning-based bimanual manipulation has gained significant attention in robotics.
-While some studies have successfully utilized demonstration data [22, 23, 24], the lack of high-quality multi-fingered hand demonstrations often restricts their applicability to simple end-effectors.
-Efforts to collect such data for multi-fingered hands have been made [25, 26, 27], but challenges such as retargeting errors and latency limit their practical use.
-On the other hand, sim-to-real approaches have also attracted growing interest in the domain of robotic manipulation [28, 29, 30, 31], demonstrating numerous successes across various tasks such as deformable object manipulation [32], open doors [33], and in-hand rotations [34].
-However, most of these methods focus on either single dexterous hands [35, 36, 37, 38] or dual-gripper systems [39, 40].
-Some progress has been made in dual-dexterous-hand manipulation [41, 42, 43].
-[41] construct a dexterous bimanual manipulation platform, offering 20 manipulation tasks.
-[42] employ multi-agent RL to train a bimanual handover task, incorporating a trajectory prediction model to bridge the sim-to-real gap.
-[43] presents a system where robot hands are trained to play the piano using DRL.
-However, these works primarily rely on visual and proprioceptive modalities, which somewhat limit the performance of manipulation tasks.
-Multimodal data has proven effective in dual-gripper [44, 20, 45] and single dexterous hand tasks [46, 47, 48], its application in bimanual dexterous manipulation remains largely unexplored.
-Our work represents a pioneering effort in applying multimodal sensing to dual-dexterous-hand manipulation.
+Although some studies have successfully used demonstration data [stepputtis2022system, zhao2023aloha, kim2021transformer], the lack of high-quality multi-fingered hand demonstrations often restricts their applicability to simple end-effectors.
+Efforts to collect such data for multi-fingered hands have been made [handa2020dexpilot, qin2022one, arunachalam2023dexterous], but challenges such as retargeting errors and motion retargeting latency limit their practical use.
+Sim-to-real approaches have demonstrated success in robotic manipulation [akkaya2019solving, liu2023dexrepnet, qin2023dexpoint, qi2023hand], including deformable object manipulation [matas2018sim], open doors [rajeswaran2017DAPG], and in-hand rotations [chen2023visual_dexterity].
+However, most methods focus on single dexterous hands [chen2023sequential, yin2023RotateWithoutSeeing, bao2023dexart, pitz2023dextrous] or dual-gripper systems [li2023efficient, lin2023bi_touch].
+Some progress has been made in dual-dexterous-hand manipulation [chen2023bi, huang2023dynamic_handover, zakka2023robopianist, Lin2024twisting][chen2023bi] construct a dexterous bimanual manipulation platform, offering 20 manipulation tasks.
+[huang2023dynamic_handover] employ multi-agent RL to train a bimanual handover task, incorporating a trajectory prediction model to bridge the sim-to-real gap.
+[zakka2023robopianist] presents a system in which robot hands are trained to play the piano using DRL.
+[Lin2024twisting] explores RL-based twisting policies for predefined cylindrical containers using sparse visual keypoints, demonstrating limited zero-shot transfer through contact-aware rewards.
+However, these works mainly rely on visual and proprioceptive modalities, which somewhat limit the performance of manipulation tasks.
+Multimodal data has proven effective in dual-gripper [hogan2020tactile, mejia2024hearing, sunil2023visuotactile_cloth] and single dexterous hand tasks [yuan2024RobotSynesthesia, qi2023inhand-visuotactile, guzey2024see], its application in bimanual dexterous manipulation remains largely unexplored.
+Our work represents a pioneering effort to apply multimodal sensing to dual-dexterous-hand manipulation.
 III Method
-In this section, we first introduce our data collection system for capturing ego-centric human manipulation data, including Vision, Tactile, Action, and Object information (VTAO), detailed in Section III-A.
-Section III-B presents VTAO-BiManip, a novel pretraining framework that utilizes human manipulation priors for multimodal feature extraction and fusion.
-Finally, Section III-C outlines the development of a bimanual dexterous bottle-cap unscrewing task in simulation and the subsequent policy training using RL.
+In this section, we first introduce our data collection system to capture ego-centric human manipulation data, which includes Vision, Tactile, Action, and Object (VTAO) information (see Section III-A).
+Section III-B then introduces VTAO-BiManip, a novel pre-training framework that utilizes human manipulation priors for multimodal feature extraction and fusion.
+Finally, Section III-C details the implementation of a dexterous bimanual bottle cap unscrewing task in simulation, followed by policy training through the curriculum RL.
 III-A VTAO Capture System for Human Bi-Manipulation
-We develop a novel VTAO data collection system for capturing multimodal data during human bimanual manipulation, built upon our previous visual-tactile acquisition system [6].
+We develop a VTAO data collection system to capture multimodal data during human bimanual manipulation.
+The visual and tactile acquisition methods are directly adopted from our previous system [liu2024m2vtp].
+The key addition is a motion capture system that records hand motion trajectories and bottle cap poses in the same coordinate system.
+Below, we detail this VTAO capture system:
 III-A1 System Overview
-The VTAO capture system includes: 1) tactile and motion acquisition gloves for both hands, 2) Hololens2  [^1] https://www.microsoft.com/en-us/hololens/buy for visual data capture, 3) the ChingMu motion capture (mocap) system  [^2] https://www.chingmu.com/ for recording hand movements and 6-DoF object pose, and 4) a personal computer responsible for data acquisition and alignment, as shown in Fig.
+The VTAO capture system includes: 1) Hololens2  [^1] https://www.microsoft.com/en-us/hololens/buy for visual data capture, 2) tactile and motion acquisition gloves for both hands, 3) the CHINGMU motion capture system  [^2] https://www.chingmu.com/ with 16 cameras to capture precise hand movements and 6-DoF object pose, and 4) a personal computer responsible for data acquisition and alignment, as shown in Fig.
 2.
-Each acquisition glove consists of two layers: an inner layer with 20 tactile sensors, as [6], and an outer layer with markers for motion capture.
-The mocap system comprises 16 cameras to capture the 6-DoF pose information of the bottle and the precise movements of the hand joints.
+The CHINGMU system plays a critical role in acquiring high-quality bimanual motion trajectories and object pose data, which are essential for the following model pre-train.
+Each acquisition glove consists of two layers: an inner layer with 20 tactile sensors (identical to the design in [liu2024m2vtp]), and an additional outer layer with retroreflective markers for optical motion capture.
 Figure 2: Our VTAO capture system for dual-hand human manipulation.
@@
 Multimodal data are captured independently, with visual data sampled at 30 Hz, tactile data at 200 Hz, and mocap data at 1 kHz.
-Data from other modalities are synchronized with the visual modality using local timestamps, which are recorded by the personal computer during the acquisition.
-To unify coordinate systems across modalities, we place markers on the Hololens2 and calibrate the transformation between the camera and the rigid body coordinates of the Hololens2.
-Utilizing this transformation allows us to map the 6-DoF pose of the bottle, as well as the hand motion data, to the RGB camera coordinate system of the visual modality, facilitating model comprehension.
+Data from other modalities are synchronized to the visual modality through local timestamps recorded by the personal computer during acquisition.
+To unify the motion capture and visual coordinate systems, we install markers on both the Hololens2 and a calibration board, then simultaneously capture their positions in the motion capture system while acquiring the calibration board’s location in the visual system, thereby calibrating the transformation between the camera coordinates and the Hololens2’s rigid-body coordinates.
+This transformation enables mapping of the 6-DoF bottle pose and hand motion trajectories to the RGB camera coordinate system of the visual modality, to facilitate model comprehension.
 III-A3 Retarget Human Manipulation Trajectories
-To unify bimanual action representations across pretraining and downstream tasks, we retarget human bimanual actions from the dataset onto the dual dexterous hands used in downstream tasks.
-Following DexRepNet [29], we regard retargeting as a nonlinear optimization problem.
-Our objective is to ensure similarity in distances from fingertips and finger segments to the wrist, in both robotic and human hands at all stages, resulting in natural robotic hand postures.
-Our objective function aligns with that presented in [25], formulated as:
+To unify bimanual action representations across pre-training and downstream tasks, we retarget human bimanual trajectories to dexterous robotic hands via nonlinear optimization, following DexRepNet [liu2023dexrepnet].
+The objective, aligned with [handa2020dexpilot], preserves geometric similarity by minimizing distances between wrist-to-fingertips and wrist-to-phalanxes across all motion phases, ensuring natural robotic hand poses, and is defined as:
 $q_{t}^{\mathbf{R}*}=\underset{q_{t}^{\mathbf{R}}}{\operatorname{argmin}}\sum_{i=0}^{N}\|\mathbf{v}^{\mathbf{R}}_{i}\left(M^{\mathbf{R}},q_{t}^{\mathbf{R}}\right)-\mathbf{v}_{i}^{\mathbf{H}}\left(M^{\mathbf{H}},q_{t}^{\mathbf{H}}\right)\|^{2},$ (1)
-where $\mathbf{v}_{i}^{\mathbf{R}}$ and $\mathbf{v}_{i}^{\mathbf{H}}$ signify the target vectors of the robotic and human hands respectively (from wrist to fingertips and wrist to finger segments).
-These are determined using forward kinematics of the dual Shadow Hand $M^{\mathbf{R}}$ and human hands $M^{\mathbf{H}}$, coupled with joint angles $q^{\mathbf{R}}_{t}$ and $q^{\mathbf{H}}_{t}$ at step $t$.
-$q_{t}^{\mathbf{R}*}$ denotes the optimized joint angles of the robotic hands at step $t$.
-We refine each step of the human bimanual manipulation trajectories within the dataset to generate retargeted robotic hand trajectories, which are subsequently utilized for the pre-training of the robots.
-Figure 3: Overview of our VTAO-BiManip pre-train framework and its downstream application.
-III-B Masked VTAO Transformer for Pretraining
-To explore the benefits of incorporating multimodal human manipulation data into dual-hand robotic skill learning, we propose a novel pre-training framework, VTAO-BiManip, based on MAE [8].
-This framework uses our collected VTAO data for pre-training.
-It extracts and integrates latent environmental representations to support the learning of downstream robotic manipulation tasks.
-VTAO-BiManip includes a VTAO encoder and a VTAO decoder, as shown in Fig.
+where $\mathbf{v}_{i}^{\mathbf{R}}$ and $\mathbf{v}_{i}^{\mathbf{H}}$ denote the target vectors for the robotic and human hands respectively, defined as wrist-to-fingertip and wrist-to-phalanx distance vectors.
+These vectors are computed via forward kinematics of the dual Shadow Hand $M^{\mathbf{R}}$ and human hand model $M^{\mathbf{H}}$, using joint angles $q^{\mathbf{R}}_{t}$ and $q^{\mathbf{H}}_{t}$ at timestep $t$.
+The optimized joint angles $q_{t}^{\mathbf{R}*}$ are obtained through this minimization process.
+We optimize each timestep of the collected human bimanual trajectories to generate retargeted robotic motion sequences, which are subsequently used for robotic policy pre-training.
+Figure 3: Overview of our VTAO-BiManip pre-train framework and its downstream bimanual manipulation with sub-skill and collaborative curriculum reinforcement learning.
+III-A4 VTAO Dataset
+Using our VTAO acquisition system, we collect 216 human bimanual manipulation trajectories with 26 different bottles, spanning 6-17 seconds per trajectory and totaling 61,684 temporally aligned multimodal frames.
+Through preprocessing, RGB images are cropped to $224\times 224$ resolution, tactile signals are binarized with a $\lambda=0.4V$ threshold, bottle poses are aligned to the camera coordinate system, and human hand motions are kinematically retargeted to dual robotic hands.
+Each processed frame includes a RGB image ($V\in\mathbb{R}^{224\times 224\times 3}$), bimanual tactile signals ($T\in\mathbb{R}^{40}$, 20 tactile sensors per hand), robotic joint angles of the current step ($A\in\mathbb{R}^{48}$, 24 joints per hand), and an object state vector ($O\in\mathbb{R}^{9}$) encoding the 6-DoF pose (position and orientation) and 3-D size information (width, height, depth) of the bottle.
+III-B Masked VTAO Transformer for Pre-Training
+To enhance dual-handed robotic skill learning through multimodal data fusion, we propose VTAO-BiManip, a MAE-based framework pre-trained on our collected VTAO dataset.
+As shown in Fig.
+3, our pre-training framework follows an encoder-decoder structure: 1) a multimodal encoder processes visual, tactile, and current action inputs for each time step from the demonstration trajectories, projecting them into latent tokens that integrate information for different modalities via attention; 2) a decoder reconstructs the original sensory modalities, predicts subsequent actions, and estimates object states from these latent tokens.
+This joint reconstruction mechanism inherently enables cross-modal feature fusion during pre-training, where the model learns to recover complete environmental interactions from partial sensory observations.
+Notably, the object state information is only used for VTAO-BiManip pre-training and is not incorporated into RL policy learning during downstream tasks to maintain practical feasibility.
+III-B1 VTAO Encoder $E_{\theta}(V,T,A)\rightarrow(h_{\text{fs}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
+The VTAO encoder processes a single frame input (V, T, A), comprising an RGB image $V$, binarized tactile signals $T$, and action $A$.
+The action $A$ is represented by robotic joint angles retargeted from human hand motions.
+Notice that the object information $O$ is excluded from inputs due to deployment constraints, but will be reconstructed via mask tokens appended to the input sequence.
+For all the input and output modalities, they are tokenized and each token is a vector of $d$ dimensions.
+The visual input ($V\in\mathbb{R}^{224\times 224\times 3}$) is divided into 196 non-overlapping patches via a Conv2D network to extract patch tokens $k_{v}\in\mathbb{R}^{196\times d}$.
+Binarized tactile signals $T\in\mathbb{R}^{40}$ (20 tactile sensors per hand) are projected into latent tokens $k_{c}\in\mathbb{R}^{40\times d}$ via an MLP, establishing a one-to-one mapping between tactile sensors and tokens.
+This binarization follows the design in [liu2024m2vtp], aiming to reduce the sim-to-real gap and eliminate the need for precise calibration across taxels.
+The action $A\in\mathbb{R}^{48}$ (joint angles from 24 joints per hand) is encoded into action tokens $k_{a}\in\mathbb{R}^{48\times d}$ through a separate MLP, with each token representing one joint angle.
+All input tokens are concatenated into a sequence $(k_{\text{fs}},k_{v},k_{c},k_{a},k_{\text{null}})$, where $k_{\text{fs}}\in\mathbb{R}^{d}$ is the fusion token (a special token aggregating global information for downstream tasks) and $k_{\text{null}}\in\mathbb{R}^{9\times d}$ are mask tokens for $O$ reconstruction.
+Following positional encoding and MAE-style masking [he2022mae], we randomly mask portions of tokens from different modalities: visual tokens $k_{v}$ (ratio $r_{v}\in[0,1)$), tactile tokens $k_{c}$ ($r_{c}\in[0,1)$), action joint tokens $k_{a}$ ($r_{a}\in[0,1)$) and object mask tokens $k_{\text{null}}$ ($r_{o}\in[0,1)$).
+The multi-modal tokens, with fusion token, are encoded by a Transformer [kim2021transformer], outputting latent representations $(h_{\text{fs}},h_{v},h_{c},h_{a},h_{o})$ for downstream tasks.
+[…]
+To enhance the feature extraction ability of the VTAO encoder $E_{\theta}$ for downstream tasks, we define four reconstruction objectives for the decoder: a) reconstruct the image $\widehat{V}$; b) reconstruct tactile signals $\widehat{T}$; c) reconstruct current actions $\widehat{A}$ and predict future actions $\widehat{A_{+1}},...,\widehat{A_{+p}}$] for $p$ steps; d) reconstruct the 6-DoF pose and 3-D size information of the bottle $\widehat{O}$, as shown in Fig.
 3 (left).
-III-B1 VTAO Encoder $E_{\theta}(V,C,A)\rightarrow(h_{\text{cls}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
-The VTA encoder takes in the visual, tactile, and action pair (V, C, A) as input.
-The RGB image $V\in\mathbb{R}^{224\times 224\times 3}$, the binarized tactile signal from both hands $C\in\mathbb{R}^{40}$ (thresholded by $\lambda$), and the joint angles $A\in\mathbb{R}^{48}$, retargeted to the dual Shadow Hand, are processed via a Conv2D and two MLPs to form a patch sequence.
-Additionally, empty patches are added to align feature dimensions.
-These patches are used by the decoder to reconstruct the bottle information $\widehat{O}$.
-After adding positional encoding, the patch sequence becomes a token sequence.
-Following MAE [8], we randomly mask $r_{v}\in[0,1)$ of the visual tokens, $r_{c}\in[0,1)$ of the tactile tokens, $r_{a}\in[0,1)$ of the action tokens, and $r_{o}\in[0,1)$ of the empty tokens to form the input token sequence $(t_{v},t_{c},t_{a},t_{\text{null}})$.
-A CLS token $t_{\text{cls}}$ is inserted into this sequence as an aggregated representation for downstream tasks.
-The final token sequence $(t_{\text{cls}},t_{v},t_{c},t_{a},t_{\text{null}})$ is processed by a Transformer [24] encoder, producing a latent representation sequence $(h_{\text{cls}},h_{v},h_{c},h_{a},h_{o})$.
-III-B2 VTAO Decoder $D_{\theta}(h_{v},h_{c},h_{a},h_{o},m)\rightarrow(\widehat{V},\widehat{C},\allowbreak[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}],\widehat{O})$
-To enhance the feature extraction ability of the VTAO encoder $E_{\theta}$ for downstream tasks, we define four reconstruction objectives for the decoder: a) reconstruct the image $\widehat{V}$; b) reconstruct tactile signals $\widehat{C}$; c) reconstruct current actions and predict future actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$; d) reconstruct the rotation and size information of the bottle $\widehat{O}$, as shown in Fig.
-3 (left).
-Here, $m$ denotes the mask tokens used for feature alignment.
-Reconstruct Image $\widehat{V}$ and Tactile Signals $\widehat{C}$
-Building on our previous work [6], we employ a Transformer Decoder and MLP layers to reconstruct the image $\widehat{V}$ and tactile signals $\widehat{C}$ from the corresponding tokens $h_{v}$, $h_{c}$, respectively.
-Serving as an auxiliary task, this approach has been demonstrated to benefit downstream applications.
-Reconstruct Current Actions and Predict Future Actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$
-To enable the encoder $E_{\theta}$ to acquire relevant prior knowledge of human manipulation tasks from single-frame VTAO input, we use a Transformer decoder and an MLP layer to reconstruct the dual-hand action of current step $\widehat{A}$ and the action sequences for the next p steps $[\widehat{A_{+1}},...,\widehat{A_{+p}}]$ from the corresponding action tokens $h_{a}$.
+Here, $m$ denotes the mask tokens used to fill in missing dimensions of the input sequence.
+Reconstruct Image $\widehat{V}$ and Tactile Signals $\widehat{T}$
+We employ a Transformer decoder and MLP layers to reconstruct the image $\widehat{V}$ and tactile signals $\widehat{T}$ from the corresponding tokens $h_{v}$ and $h_{c}$, respectively.
+This approach has been demonstrated to benefit downstream applications in [liu2024m2vtp].
+Reconstruct Current Actions and Predict Future Actions $\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}$]
+To enable the encoder $E_{\theta}$ to acquire prior knowledge of human manipulation tasks from single-frame VTAO input, we use a Transformer decoder and an MLP layer to reconstruct the current dual-hand action $\widehat{A}$ and predict future actions $\widehat{A_{+1}},...,\widehat{A_{+p}}$] for p steps, based on the action tokens $h_{a}$.
 Reconstruct Bottle Informations $\widehat{O}$
-Knowing the pose and size of the object is beneficial for downstream tasks such as bottle cap twisting.
-We enable the network to reconstruct the 6-DoF pose of the bottle relative to the camera coordinate system, along with the dimensions of the bottle cap and body, from object tokens $h_{\text{o}}$.
-This enables the encoder $E_{\theta}$ to better focus on the relevant information of the manipulated object.
+Our framework reconstructs the 6-DoF pose of the assembled object (bottle with an attached cap) relative to the camera’s 3D coordinate frame, along with its 3D bounding box dimensions (width, height, depth).
+This is achieved by decoding the object tokens $h_{\text{o}}$, which forces the encoder $E_{\theta}$ to learn geometry-dynamic representations.
+These representations encode structural priors of the object and its spatiotemporal state transitions, thereby enabling real-time motion adaptation during dynamic manipulation.
 Loss Design
 We formulate the loss function as:
-$L(\theta)=$ $W_{\text{img}}\cdot L(V,\widehat{V})+W_{\text{tac}}\cdot L(C,\widehat{C})+W_{\text{bot}}\cdot L(O,\widehat{O})$
-$W_{\text{act}}\cdot L([A,A_{+1},...,A_{+p}],[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]),$ (2)
-where $W_{\text{img}},W_{\text{tac}},W_{\text{bot}}$, and $W_{\text{act}}$ are weights and set to 1, 2, 5, 2, respectively. $L(\cdot)$ denotes the Euclidean distance between two vectors.
-III-C VTAO RL for bimanual manipulations
-III-C1 Problem Formulation
+[…]
+[…]
+where $W_{\text{img}},W_{\text{tac}},W_{\text{bot}}$, and $W_{\text{act}}$ are weights for each loss term, set to 1, 2, 5, and 2, respectively.
+$L(\cdot)$ denotes the Euclidean distance between two vectors.
+III-C Curriculum RL for Dexterous Bimanual Manipulation
+III-C1 Task
+To assess the effectiveness of VTAO pre-training on downstream tasks, we implement a bimanual bottle-cap rotation task (focusing solely on the unscrewing process).
+The setup positions the bottle body at the center of the table with the left hand in a pre-grasp position (left side) and the right hand positioned above it, as shown in the submitted video.
+The bottle body and cap remain connected through a z-axis rotational joint.
+Successful operation is defined by cap rotation exceeding 4 rad (angular displacement) rather than physical removal, as defined in [liu2024m2vtp].
+III-C2 Problem Formulation
 We model the bimanual manipulation task as a Markov Decision Process (MDP) defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$.
@@
 3 (right).
-III-C2 Task
-To assess the effectiveness of VTAO pre-training on downstream tasks, we implement a bimanual bottle-cap unscrewing task in Isaac gym [49].
-The setup positions the bottle body at the center of the table, with the left hand to its left in a pre-grasp gesture and the right hand fixed above it, as shown in the submitted video.
-The bottle body and cap are linked by a joint that rotates around the z-axis.
-III-C3 State Space $\mathcal{S}$
-We define the state space as $\mathcal{S}=\{V_{\text{obs}},C_{\text{obs}},\mathcal{P}\}$, where $V_{\text{obs}}$ is the ego-centric RGB image, $C_{\text{obs}}$ is the binary tactile signal from both hands, thresholded by $\lambda$, and $\mathcal{P}$ is proprioception data of both hands, including joint angles and velocities.
+III-C3 State Space with Multi-Modal Pre-trained Representations $\mathcal{S}$
+We define the state space as $\mathcal{S}=\{V_{\text{obs}},T_{\text{obs}},\mathcal{P}\}$, where $V_{\text{obs}}$ is the ego-centric RGB image, $T_{\text{obs}}$ is the binary tactile signal from both hands, thresholded by $\lambda$, and $\mathcal{P}$ is proprioception data of both hands, including joint angles and velocities.
+Notably, object state information is excluded from the input state space $\mathcal{S}$ during RL training, as such privileged data is difficult to directly measure under real-world deployment constraints.
 To maintain feature representation from pre-training and improve stability in RL with sparse rewards, we freeze the parameters of the pre-trained VTAO model during RL training.
-The final feature vector input to the policy network is $\{h_{\text{cls}},\phi(\mathcal{P})\}$, where $h_{\text{cls}}=\mathrm{E}_{\theta_{f}}(V_{\text{obs}},C_{\text{obs}},A_{\text{obs}})$, with $\mathrm{E}_{\theta_{f}}$ representing the pre-trained VTAO encoder with frozen parameters.
+The final feature vector input to the policy network is $\{h_{\text{fs}},\phi(\mathcal{P})\}$, where $h_{\text{fs}}=\mathrm{E}_{\theta_{f}}(V_{\text{obs}},T_{\text{obs}},A_{\text{obs}})$, with $\mathrm{E}_{\theta_{f}}$ representing the pre-trained VTAO encoder with frozen parameters.
 $A_{\text{obs}}\in\mathcal{P}$ represents the joint angles of both hands.
-$h_{\text{cls}}$ is the CLS token output by $\mathrm{E}_{\theta_{f}}$.
+$h_{\text{fs}}$ is the FS token output by $\mathrm{E}_{\theta_{f}}$.
 $\phi(\cdot)$ is a linear layer.
 III-C4 Action Space $\mathcal{A}$
-In the simulation, we use Shadow Hand [1] as the manipulator.
+In the simulation, we use Shadow Hand [shadowhand] as the manipulator.
 This five-fingered robotic hand features 24 DoFs.
 Excluding the thumb, the distal joints of the four fingers are tendon-driven, resulting in a total of 20 DoFs.
-We implement dual-hand manipulation, where the base link of the right hand is fixed and the left hand can move freely with six DoFs.
+We implement dual-hand manipulation, where the right hand is controlled with its initial position serving as the reference point, while the left hand is freely manipulated with six DoFs.
 Consequently, the action $a=\pi_{\theta}(s)\in\mathbb{R}^{46}$.
-III-C5 Learning Curriculum
+III-C5 Learning Curriculum with Stage-specific Reward Design $\mathcal{R}$
 It is challenging to train precise dual-hand manipulation with non-fixed objects from scratch.
-To address this, we use a two-stage learning curriculum to stabilize the learning process.
+In our experiments, single-stage training yields a consistent zero success rate due to insufficient exploration in the high-dimensional bimanual action space.
+To overcome this and stabilize the learning process, we implement a progressive two-stage curriculum with phase-dependent reward functions $r_{n}(n\in\{1,2\})$, where each stage’s total reward combines the individual rewards of the left and right hands: $r_{n}=r_{\text{left}_{n}}+r_{\text{right}_{n}}$.
+Stage 1: Sub-skill Learning.
 In the first stage, we fix the bottle body on the table, enabling the left hand to learn to grasp the bottle and the right hand to learn to unscrew the cap.
-In the second stage, we release the bottle body, allowing both hands to collaborate in unscrewing the cap.
-III-C6 Reward Function $\mathcal{R}$
-To coordinate the dual dexterous hands in complex action spaces across different training stages, our reward function $r_{n}(n\in\{1,2\})$ is divided into two phases.
-In each phase, the total reward is the sum of the individual rewards for the left and right hands, $r_{n}=r_{\text{left}_{n}}+r_{\text{right}_{n}}$.
 The reward function of the first stage $r_{1}$ is:
-{$r_{\text{left}_{1}}&=r_{\text{hdis}}+r_{\text{fcon}}\\ r_{\text{right}_{1}}&=r_{\text{cang}}+r_{\text{cvel}}+r_{\text{fdis}}\end{aligned}\right..$ (3)
-Here, $r_{\text{hdis}}=\alpha_{1}\cdot d_{\text{h2b}}$ is a distance penalty, where $d_{\text{h2b}}$ is the distance between the left hand and the bottle.
-$r_{\text{fcon}}=\alpha_{2}\cdot N_{\text{con}}$ incentivizes multiple contacts between the left hand and the bottle, with $N_{\text{con}}$ representing the number of contacts.
-The rotation angle reward, $r_{\text{cang}}=\alpha_{3}\cdot\min(a_{\text{c}},7.0)$, depends on the rotation angle of the bottle cap $a_{\text{c}}$.
-$r_{\text{cvel}}=\alpha_{4}\cdot C_{\text{flag}}\cdot v_{\text{c}}$ rewards rotation speed, where $v_{\text{c}}$ is the rotation speed of the cap, and $C_{\text{flag}}$ equals 1 if the right hand is in contact with the cap, and 0 otherwise.
-The term $r_{\text{fdis}}=\alpha_{5}\cdot\exp(-10\cdot d_{\text{fz}})$ encourages the right-hand fingertips to approach the bottle cap, with $d_{\text{fz}}$ being the z-axis distance between the fingertips and the height of the cap.
-The weights $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{5}$ are set to $-5$, $0.05$, $0.5$, $1.1$, and $0.5$, respectively.
-The rewards for the second phase, $r_{2}$ are as follows:
-{$r_{\text{left}_{2}}&=r_{\text{htdis}}+r_{\text{bdis}}+r_{\text{brot}}\\ r_{\text{right}_{2}}&=r_{\text{right}_{1}}\end{aligned}\right.,$ (4)
-where $r_{\text{htdis}}=\beta_{1}\cdot\exp(-5\cdot d_{\text{h2t}})$ is the approaching reward, with $d_{\text{h2t}}$ representing the distance between the left hand and the target position.
-$r_{\text{bdis}}=\beta_{2}\cdot\exp(-10\cdot d_{\text{o2i}})$ encourages the bottle to remain near its initial position, with $d_{\text{o2i}}$ denoting the distance between the bottle’s current and initial position.
-The reward $r_{\text{brot}}=\beta_{3}\left(\frac{1}{|d_{\text{qua}}|+1}\right)$ encourages the axis of the bottle remains perpendicular to the table, where $d_{\text{qua}}=2\cdot\arcsin\left(\min\left(\lVert q_{\text{diff}}\rVert_{2},1.0\right)\right)$ , and $q_{\text{diff}}=q_{\text{bot}}\cdot q_{\text{ini}}$ represents the vector part of the quaternion difference between the bottle’s current quaternion $q_{\text{bot}}$ and its initial orientation $q_{\text{ini}}$.
-The weights $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are all set to 1.0.
+  • Left hand learns bottle grasping through:
+    $r_{\text{left}_{1}}=\underbrace{\alpha_{1}d_{\text{h2b}}}_{\text{distance penalty}}+\underbrace{\alpha_{2}N_{\text{con}}}_{\text{contact reward}}.$ (3)
+    Here, $\alpha_{1}\!=\!-5$, $\alpha_{2}\!=\!0.05$.
+$d_{\text{h2b}}$ is the distance between the left hand and the bottle.
+$N_{\text{con}}$ represents the number of contacts between the left hand and the bottle body.
+  • Right hand develops cap rotation skills via:
+    $r_{\text{right}_{1}}=\underbrace{\alpha_{3}\min(a_{c},4.0)}_{\text{rotation angle reward}}+\underbrace{\alpha_{4}C_{\text{flag}}v_{c}}_{\text{rotation speed reward}}+\underbrace{\alpha_{5}e^{-10d_{\text{fz}}}}_{\text{approach reward}}.$ (4)
+    Here, $\alpha_{3}\!=\!0.5$, $\alpha_{4}\!=\!1.1$, $\alpha_{5}\!=\!0.5$.
+$a_{\text{c}}$ is the rotation angle of the bottle cap.
+$C_{\text{flag}}$ equals 1 if the right hand is in contact with the cap, and 0 otherwise.
+$v_{\text{c}}$ is the rotation speed of the cap.
+The approach reward term encourages the right-hand fingertips to approach the bottle cap, with $d_{\text{fz}}$ denoting the z-axis distance between the fingertips and cap’s top surface.
+Stage 2: Collaborative Learning. In the second stage, we release the bottle body, allowing both hands to collaborate in unscrewing the cap. The rewards for the second stage, $r_{2}$ are as follows:
+  • Left hand manages grasping and stabilization:
+    $r_{\text{left}_{2}}=\underbrace{\beta_{1}e^{-5d_{\text{h2t}}}}_{\text{hand approach}}+\underbrace{\beta_{2}e^{-10d_{\text{o2i}}}}_{\text{bottle stability}}+\underbrace{\frac{\beta_{3}}{|d_{\text{qua}}|+1}}_{\text{upright maintenance}}.$ (5)
+    Here, $\beta_{1}\!=\!\beta_{2}\!=\!\beta_{3}\!=\!1.0$.
+$d_{\text{h2t}}$ is the distance between the left hand and the target position $p_{t}=(x_{0},y_{0}-0.06,z_{0})$, with $(x_{0},y_{0},z_{0})$ being the hand’s initial position.
+The bottle stability term encourages the bottle to remain near its initial position, with $d_{\text{o2i}}$ denoting the distance between the bottle’s current and initial position.
+The upright maintenance term encourages the axis of the bottle remains perpendicular to the table, where $d_{\text{qua}}=2\cdot\arcsin\left(\min\left(\lVert q_{\text{diff}}\rVert_{2},1.0\right)\right)$, and $q_{\text{diff}}=q_{\text{bot}}\cdot q_{\text{ini}}$ represents the vector part of the quaternion difference between the bottle’s current quaternion $q_{\text{bot}}$ and its initial orientation $q_{\text{ini}}$.
+  • Right hand maintains stage 1 rewards $r_{\text{right}_{2}}=r_{\text{right}_{1}}$ while adapting to dynamic bottle movements.
 IV Experiment Results
@@
 IV-A Experiment Settings
-IV-A1 Dataset
-Using our VTAO acquisition system, we collect 216 human bimanual manipulation trajectories involving 26 different bottles.
-Each trajectory lasts 6 to 17 seconds and comprises a total of 61684 frames.
-Each frame contains aligned visual, tactile, and action data, along with bottle pose and size labels, as shown in the submitted video.
-Before training VTAO-BiManip, we preprocess the data by cropping images to $224\times 224$ pixels, binarizing tactile data with a $\lambda=0.4V$ threshold, aligning bottle pose with the camera coordinate system of the visual data, and retargeting human hand movements to the joint angles of the dual dexterous hands.
-The final data representations include images $V\in\mathbb{R}^{224\times 224\times 3}$, tactile data $C\in\mathbb{R}^{40}$, and action data $A\in\mathbb{R}^{48}$.
-IV-A2 Metrics
+IV-A1 Metrics
 We use success rate to evaluate the effectiveness of our method on downstream tasks.
-A task is considered successful if the dexterous hand rotates the bottle cap by more than half a turn.
-During training, we calculate the success rate as the average of the last ten manipulations in each parallel environment.
-For evaluation, we select the model from the 3500th iteration of the second training phase and measure the average success rate across 10 manipulations for each bottle in the evaluation set.
-IV-A3 Implementation Details
-We select 15 bottles from ShapeNet [50] with varying body and cap sizes, using 10 for training (Seen) and 5 for testing (Unseen).
+A task is considered successful only if the dexterous hand rotates the bottle cap by more than 4 rad.
+During training, we compute the success rate for each iteration as the moving average over the preceding ten episodes across all parallel environments.
+For evaluation, we select the model from the 3500th iteration of the second training stage and measure the average success rate across 10 manipulations for each bottle in the evaluation set.
+IV-A2 Implementation Details
+We deploy our bimanual bottle cap twisting task in the Isaac Gym simulation environment [makoviychuk2021isaac].
+We select 15 bottles from ShapeNet [chang2015shapenet] with varying body and cap sizes, using 10 for training (Seen) and 5 for testing (Unseen).
 Our VTAO-BiManip runs on a system with an Intel Xeon Gold 6326 and an NVIDIA 3090.
@@
 Training with a minibatch size of 8 takes about 2 hours for 300 epochs.
-For RL, we run 400 parallel environments and train using the PPO algorithm for roughly 4500 iterations, requiring about 62 hours to train the dual-hand manipulation strategy network.
-IV-A4 Baselines
-To validate the effectiveness of our method, we design several baselines with different pre-training structures, as shown in Table I.
-We designate our baselines in the form of VTAO.
-Here, V, T, and A correspond to the pretraining of visual, tactile, and action modalities, respectively, while O represents the reconstruction of bottle information.
-Both tactile and action modalities involve data from two hands.
-The action modality entails reconstructing the current frame and predicting the next five frames.
-The absence of a letter in the naming indicates that the corresponding modality is not used.
-For methods that do not involve pretraining, we denote them with the suffix ‘-Scr’.
+For RL, the two-stage curriculum learning takes approximately 62 hours to learn the dual-hand manipulation strategy.
+In the first stage, we train the policy using PPO for 1000 iterations.
+The resulting model is then used as the initialization for the second stage, where we further train for 3500 iterations.
+IV-A3 Baselines
+To validate the effectiveness of our method, we design multiple baseline models with different configurations, as summarized in Table 1.
+The baselines follow a naming convention of VTAO, where each capital letter indicates the inclusion of a specific pretrained modality using the MAE architecture shown in Figure 3: V-Visual (RGB observations), T-Bimanual binary tactile signals, A-Bimanual action joints (current reconstruction + 5-frame prediction) and O - Bottle information reconstruction.
+The absence of a letter denotes exclusion of the corresponding modality.
+We also construct a baseline that trains the multi-modal representation and the action policy from scratch by end-to-end reinforcement learning VTA-Scr.
+The baseline employs a ResNet encoder for visual inputs and MLP networks for processing tactile and action modalities.
 TABLE I: Settings for baselines.
 Methods Modality Pretrain JointPretrain ReconBottle
+VTA-Scr V+T+A $\times$ $\times$ $\times$
 V V $\checkmark$ $\times$ $\times$
@@
 VTO V+T $\checkmark$ $\checkmark$ $\checkmark$
-VTA-Scr v+t+a $\times$ $\times$ $\times$
 VTAO (Ours) V+T+A $\checkmark$ $\checkmark$ $\checkmark$
@@
 * indicates that the method failed to train.
-Figure 4: Qualitative results of different pre-train methods. The left is the training process; the right is the evaluation results
+Figure 4: Qualitative results of different pre-train methods in simulation. Left: training process; Right: evaluation results.
 We design experiments to validate the effectiveness of our proposed multimodal joint pre-training. The experimental results are illustrated in the Fig. 4.
@@
 This highlights the effectiveness of our multimodal joint pretraining strategy.
-By comparing VT, VTA, VTO, and VTAO, we find that incorporating action modality prediction and object understanding into visuo-tactile fusion training enhances downstream task success rates.
-This validates the effectiveness and complementarity of the action prediction task and the object understanding task proposed in our approach.
-Both Base and A, which only possess proprioception of the robotic hands without environmental perception, fail in training.
-In contrast, the pretraining method using only tactile perception, T, demonstrates a certain level of manipulation capability, effectively demonstrating the ability of tactile perception to gather environmental information and its effectiveness in enhancing operational performance.
+Analysis of modality integration reveals two key findings: First, the performance improvements from VT to VTA (incorporating action modality prediction) and to VTO (incorporating object understanding) demonstrate that both action prediction and object understanding significantly enhance visuo-tactile fusion.
+Furthermore, VTAO achieves the highest success rates by combining both modalities, indicating complementary benefits between action prediction and object understanding tasks.
+Experiments with proprioception-only configurations (Base and A) fail to train, whereas Base+Obj achieves successful training.
+This underscores the critical importance of object state perception for enabling policy networks to dynamically adjust manipulation strategies based on target object status.
+Although our method achieves the highest success rate, the overall performance (around 60%) remains relatively low, which we attribute to the task’s high complexity — specifically, the large variations in object shapes and sizes in our experimental setting and the high-dimensional action space of dual-hand manipulation.
 IV-C Ablation Study
+We conduct comprehensive ablation studies to validate key design choices in our approach. All experimental results are summarized in the Fig. 5.
 IV-C1 Effectiveness of Bimanual Pre-training
@@
 Methods Tac Act PredictAct ActToken Obj Seen Unseen
-Ours dual dual 5steps $24\times 2$checkmark 63±4 51±7
+Ours dual dual 5steps $24\times 2$ $\checkmark$ 63±4 51±7
 v1 dual dual 5steps $24\times 2$ $\times$ 54±23 39±15
@@
 Figure 6: Our manipulation platform.
-We create a real-world experimental platform with dual Leap Hands [2] and a Unitree Z1 arm  [^3] https://www.unitree.com/z1, as shown in Fig.
+We create a real-world experimental platform with dual Leap Hands [shaw2023leaphand] equipped with tactile sensors and a Unitree Z1 arm  [^3] https://www.unitree.com/z1, as shown in Fig.
 6.
-The simulation task for bottle cap unscrewing is recreated in Isaac Gym [49] using dual Leap Hands.
+Visual perception is obtained through a ego-centric RGB camera, while tactile signals is captured via tactile sensors mounted on the dual Leap Hands and processed through binarization.
+The simulation task for bottle cap unscrewing is recreated in Isaac Gym [makoviychuk2021isaac] using dual Leap Hands.
 The final validation confirms that our method supports effective physical experimentation.
+Please refer to our submitted video for detailed demonstrations of the system’s real-world performance.
 V Conclusion
-This paper investigates the effectiveness of a pre-trained representation model integrating multi-modalities, action prediction, and object understanding for dexterous bimanual manipulation.
+This paper investigates the effectiveness of a pre-trained representation model integrating multi-modal input, action prediction, and object understanding for dexterous bimanual manipulation.
 We design a novel pre-training framework, VTAO-BiManip, which is pre-trained using human VTAO demonstrations.
@@
 This study represents a pioneering effort in multimodal bimanual manipulation.
-Future work may focus on: 1) developing large-scale multimodal bimanual datasets across diverse scenes and tasks; 2) exploring additional bimanual manipulation tasks; 3) leveraging unpaired existing 3D object labeling data.
+Future work may focus on: 1) Scaling multimodal bimanual datasets across diverse scenes and tasks; 2) Extending to broader manipulation scenarios; 3) Exploring geometric object priors from vision models.
+Our method utilizes motion-captured object 3D bounding boxes (pose+size) for the object understanding.
+However, it may be replaced by poses estimated by large models for object pose estimaiton in the future.